I have two aquariums. One is 12x18x30. The other is 13x24x30. how many gallons of water will fit into each of them.
Answer by mananth(16949)

I presume they are in feet

One is 12x18x30.
Volume = L*W*H
12x18x30=6480 cu.ft


The other is 13x24x30
=9360
9360+6480 =14840 cu.ft
1 cu.ft = 7.48 gallons
14840/7.48= 1984 gallons ( rounded off)